# Simulating and probing many-body quantum states in waveguide-QED systems with giant atoms

### v1

@@ III.5 Influence of spontaneous emission to channels other than the waveguide continuum
In practice, the unavoidable photon loss resulting from the coupling of the system with the surrounding environmental degrees of freedom should be taken into account. Assuming that all atoms have the same photon loss rate $\gamma_{0}$, the Hamiltonian (1) of the system can be rewritten as the following non-Hermitian one $\hat{H}^{\prime}=\hat{H}-(\mathrm{i}\gamma_{0}/2)\sum_{i=1}^{N}{\hat{\sigma}_{i}^{+}\!\>\hat{\sigma}_{i}^{-}}$. The corresponding scattering amplitudes can be obtained by replacing the term $\mathbf{H}$ in Eqs. (6a) and (6b) by $\mathbf{H}-{\mathrm{i}\gamma_{0}}\mathbf{I}/2$. And the effective decay of the $n$th collective mode becomes $\tilde{\Gamma}_{n}+\gamma_{0}$. Obviously, the collective modes with small coupling rates to the waveguide (with $\tilde{\Gamma}_{n}\ll\gamma_{0}$) are hardly excited due to their extremely low coupling efficiency to the waveguide. Thus it can be observed from the transmission spectrum that the dips with narrower widths (corresponding to the subradiant states) are more significantly suppressed when a photon-loss rate $\gamma_{0}$ is included, as shown in Fig. 7. For an atom array with dissipation rate of $\gamma_{0}=10^{-3}\gamma$ (which can be achieved in current wQED setups based on superconducting quantum circuits Mirhosseini et al. (2019)), all the collective modes of the atom array can still be probed by the transmission spectrum, including the extremely subradiant ones near the band edges, as shown by the blue dashed line in Fig. 7.
@@
In summary, we present an experimentally feasible scheme to simulate and probe the AAH model by utilizing giant-atom wQED systems. In our proposal, the coupling between a pair of neighboring atoms is generated by the nearly decoherence-free interaction mediated by the waveguide modes, thus eliminating the need for direct couplings between atoms. In addition to mediating interactions between atoms, the photonic modes in the waveguide are also used to detect the energy spectra of the atom array. Moreover, our method allows the energy spectrum of the atomic chain to be revealed by a single scattering process. As a result, the proposed system is characterized by a simple structure and high detection efficiency. In addition, the optimized parameters for achieving high-precision quantum simulation are analyzed to demonstrate the feasibility of the scheme under existing experimental techniques.
@@ Appendix A Scattering amplitudes expressed in terms of collective modes
$t=1-\mathrm{i}\sum_{n=1}^{N}\frac{\big{(}\mathbf{V}^{\dagger}{\mathbf{U}^{\mathrm{\mathscr{R}}}_{n}}\big{)}\big{(}{\mathbf{U}^{\mathscr{L}}_{n}}^{\dagger}\mathbf{V}\big{)}}{\Delta-\lambda_{n}},$ (19a)
@@
where $\mathbf{U}^{\mathscr{R}}_{n}$ and $\mathbf{U}^{\mathscr{L}}_{n}$ are the right and left eigenvectors of the non-Hermitian Hamilton matrix $\mathbf{H}$, and $\lambda_{n}$ is the corresponding complex eigenvalues, satisfying $\mathbf{H}\mathbf{U}^{\mathscr{R}}_{n}=\lambda_{n}\mathbf{U}^{\mathscr{R}}_{n}$, $\mathbf{H}^{{\dagger}}\mathbf{U}^{\mathscr{L}}_{n}=\lambda^{*}_{n}\mathbf{U}^{\mathscr{L}}_{n}$, and ${\mathbf{U}^{\mathscr{L}}_{n}}^{\dagger}\mathbf{U}^{\mathscr{R}}_{n^{\prime}}={\mathbf{U}^{\mathscr{R}}_{n}}^{\dagger}\mathbf{U}^{\mathscr{L}}_{n^{\prime}}=\delta_{nn^{\prime}}$ Brody (2013). The numerators in Eqs. (19a) and (19b) represent the overlap degree of the $n$th collective state of atomic chain and the propagating photon modes.
Equations (19a) and (19b) show that the scattering spectrum can be regarded as the result of interference between different scattering channels, which are provided by the corresponding collective modes. Under the Markovian approximation, the input $\mathbf{V}$, the Hamiltonian $\mathbf{H}$, and consequently the eigenvalues $\lambda_{n}$ and eigenvectors $\mathbf{U}^{\mathscr{R,L}}_{n}$ of $\mathbf{H}$, are independent of the frequency of the photons. Accordingly, the eigen frequency (with respect to the reference frequency $\omega_{\mathrm{a}}$) and the effective decay of the $n$th collective mode can be defined as $\tilde{\Delta}_{n}=\mathrm{Re}(\lambda_{n})$ and $\tilde{\Gamma}_{n}=-2\mathrm{Im}(\lambda_{n})$, respectively. Further, we can define the weight factors $\eta_{n}=-2\mathrm{i}[{\big{(}\mathbf{V}^{\dagger}{\mathbf{U}^{\mathscr{R}}_{n}}\big{)}\big{(}{\mathbf{U}^{\mathscr{L}}_{n}}^{\dagger}\mathbf{V}\big{)}}]/\tilde{\Gamma}_{n}$ and $\xi_{n}=-2\mathrm{i}[{\big{(}\mathbf{V}^{\top}{\mathbf{U}^{\mathscr{R}}_{n}}\big{)}\big{(}{\mathbf{U}^{\mathscr{L}}_{n}}^{\dagger}\mathbf{V}\big{)}}]/\tilde{\Gamma}_{n}$. Thus the transmission and reflection amplitudes can be written as

### v2

@@ III.5 Influence of spontaneous emission to channels other than the waveguide continuum
In practice, the unavoidable photon loss resulting from the coupling of the system with the surrounding environmental degrees of freedom should be taken into account. Assuming that all atoms have the same photon loss rate $\gamma_{0}$, the Hamiltonian (1) of the system can be rewritten as the following non-Hermitian one $\hat{H}^{\prime}=\hat{H}-(\mathrm{i}\gamma_{0}/2)\sum_{i=1}^{N}{\hat{\sigma}_{i}^{+}\!\>\hat{\sigma}_{i}^{-}}$. The corresponding scattering amplitudes can be obtained by replacing the term $\mathbf{H}$ in Eqs. (6a) and (6b) by $\mathbf{H}-{\mathrm{i}\gamma_{0}}\mathbf{I}/2$. And the effective decay of the $n$th collective mode becomes $\tilde{\Gamma}^{\prime}_{n}=\tilde{\Gamma}_{n}+\gamma_{0}$. Obviously, the collective modes with small coupling rates to the waveguide (with $\tilde{\Gamma}_{n}\ll\gamma_{0}$) are hardly excited due to their extremely low coupling efficiency to the waveguide. Thus it can be observed from the transmission spectrum that the dips with narrower widths (corresponding to the subradiant states) are more significantly suppressed when a photon-loss rate $\gamma_{0}$ is included, as shown in Fig. 7. For an atom array with dissipation rate of $\gamma_{0}=10^{-3}\gamma$ (which can be achieved in current wQED setups based on superconducting quantum circuits [73]), all the collective modes of the atom array can still be probed by the transmission spectrum, including the extremely subradiant ones near the band edges, as shown by the blue dashed line in Fig. 7.
Now we discuss the influence of photon loss on probing the many-body localization transition. When a photon loss $\gamma_{0}\ll\Gamma_{\text{eff}}$ is included, all the line widths (denote as $\tilde{\Gamma}^{\prime}_{n}=\tilde{\Gamma}_{n}+\gamma_{0}$), as well as their average (denote as $\Gamma^{\prime}_{\text{eff}}=\sum_{n}\tilde{\Gamma}’_{n}/N=\Gamma_{\text{eff}}+\gamma_{0}$, which is equal to the total decay rate of a single atom), undergo an identical increment $\gamma_{0}$. Thus, the width variance remains unchanged even after accounting for photon loss, i.e., $\sigma^{\prime 2}=\sum_{n}(\tilde{\Gamma}^{\prime}_{n}-\Gamma^{\prime}_{\text{eff}})^{2}/N=\sigma^{2}$ [see Eq. (19)] is fulfilled. Therefore, as long as $\gamma_{0}\ll\Gamma_{\text{eff}}$ is satisfied so that all the widths $\tilde{\Gamma}^{\prime}_{n}$ of the resonance dips can be measured, the corresponding variance $\sigma^{\prime 2}$ can still accurately indicate whether the system is in the localized or extended phase. On the other hand, using these measured $\tilde{\Gamma}^{\prime}_{n}$, the IPR describing the distribution of the decay rates $\text{IPR}^{\prime}_{\text{decay}}=\sum_{n}\tilde{\Gamma}^{\prime 2}_{\text{norm},n}$ can also be constructed [similar to $\text{IPR}_{\text{decay}}$ defined in Eq (20)], with $\tilde{\Gamma}^{\prime}_{\text{norm},n}=\tilde{\Gamma}^{\prime}_{n}/(N\Gamma^{\prime}_{\text{eff}})$ and $\sum_{n}\tilde{\Gamma}^{\prime}_{\text{norm},n}=1$. It can be easily proved that in the thermodynamic limit, the relation $\text{IPR}^{\prime}_{\text{decay}}\simeq(1-2\gamma_{0}/\Gamma_{\text{eff}})\text{IPR}_{\text{decay}}$ is satisfied. Clearly, $\text{IPR}^{\prime}\simeq 0$ ($\text{IPR}^{\prime}>0$) is satisfied when the system is in the localized (extended) phase. Therefore, the quantity $\text{IPR}^{\prime}_{\text{decay}}$ can also characterize the many-body quantum phase transition in the photon-loss case.
@@
In summary, we present an experimentally feasible scheme to simulate and probe the AAH model by utilizing giant-atom wQED systems. In our proposal, the coupling between a pair of neighboring atoms is generated by the nearly decoherence-free interaction mediated by the waveguide modes, thus eliminating the need for direct couplings between atoms. In addition to mediating interactions between atoms, the photonic modes in the waveguide are also used to detect the energy spectrum of the atom array. Moreover, our method allows the energy spectrum of the atomic chain to be revealed by a single scattering process. As a result, the proposed system is characterized by a simple structure and high detection efficiency. In addition, the optimized parameters for achieving high-precision quantum simulation are analyzed to demonstrate the feasibility of the scheme under existing experimental techniques.
@@ Appendix A Scattering amplitudes expressed in terms of collective modes
$t=1-\mathrm{i}\sum_{n=1}^{N}\frac{\big{(}\mathbf{V}^{\dagger}{\mathbf{U}^{\mathscr{R}}_{n}}\big{)}\big{(}{\mathbf{U}^{\mathscr{L}}_{n}}^{\dagger}\mathbf{V}\big{)}}{\Delta-\lambda_{n}},$ (21a)
@@
where $\mathbf{U}^{\mathscr{R}}_{n}$ and $\mathbf{U}^{\mathscr{L}}_{n}$ are the right and left eigenvectors of the non-Hermitian Hamilton matrix $\mathbf{H}$, and $\lambda_{n}$ is the corresponding complex eigenvalues, satisfying $\mathbf{H}\mathbf{U}^{\mathscr{R}}_{n}=\lambda_{n}\mathbf{U}^{\mathscr{R}}_{n}$, $\mathbf{H}^{{\dagger}}\mathbf{U}^{\mathscr{L}}_{n}=\lambda^{*}_{n}\mathbf{U}^{\mathscr{L}}_{n}$, and ${\mathbf{U}^{\mathscr{L}}_{n}}^{\dagger}\mathbf{U}^{\mathscr{R}}_{n^{\prime}}={\mathbf{U}^{\mathscr{R}}_{n}}^{\dagger}\mathbf{U}^{\mathscr{L}}_{n^{\prime}}=\delta_{nn^{\prime}}$ [74]. The numerators in Eqs. (21a) and (21b) represent the overlap degree of the $n$th collective state of atomic chain and the propagating photon modes.
Equations (21a) and (21b) show that the scattering spectrum can be regarded as the result of interference between different scattering channels, which are provided by the corresponding collective modes. Under the Markovian approximation, the input $\mathbf{V}$, the Hamiltonian $\mathbf{H}$, and consequently the eigenvalues $\lambda_{n}$ and eigenvectors $\mathbf{U}^{\mathscr{R,L}}_{n}$ of $\mathbf{H}$, are independent of the frequency of the photons. Accordingly, the eigen frequency (with respect to the reference frequency $\omega_{\text{a}}$) and the effective decay of the $n$th collective mode can be defined as $\tilde{\Delta}_{n}=\mathrm{Re}(\lambda_{n})$ and $\tilde{\Gamma}_{n}=-2\mathrm{Im}(\lambda_{n})$, respectively. Further, we can define the weight factors $\eta_{n}=-2\mathrm{i}[{\big{(}\mathbf{V}^{\dagger}{\mathbf{U}^{\mathscr{R}}_{n}}\big{)}\big{(}{\mathbf{U}^{\mathscr{L}}_{n}}^{\dagger}\mathbf{V}\big{)}}]/\tilde{\Gamma}_{n}$ and $\xi_{n}=-2\mathrm{i}[{\big{(}\mathbf{V}^{\top}{\mathbf{U}^{\mathscr{R}}_{n}}\big{)}\big{(}{\mathbf{U}^{\mathscr{L}}_{n}}^{\dagger}\mathbf{V}\big{)}}]/\tilde{\Gamma}_{n}$. Thus the transmission and reflection amplitudes can be written as